IF triangle abc is a right triangle and AB=x and Bc=x+1 and Ac=x+9 what are all the values of x?
Answer by shree840(260)

Right angle triangle ABC
ac^2=ab^2+bc^2
or (x+9)^2=x^2+(x+1)^2
or x^2+81+18x=x^2+x^2+1+2x
or 80= X^2-16x
or x^2-16x-80=0
x^2-20x+4x-80=0
x(x-20)+4(x-20)=0
(x+4)(x-20)=0
either x=-4 or x=20
since x is a length it can not be negative so x=20 ANS